A rectangular box measuring 150cm by 156cm by 216cm is to be packed with exact number of cubes of maximum lengths. Calculate then volume of one such cube
Answer by ikleyn(52781)

.
A rectangular box measuring 150cm by 156cm by 216cm is to be packed with exact number
of cubes of maximum lengths. Calculate then volume of one such cube
~~~~~~~~~~~~~~~~

This problem is to find Greatest Common Divisor (GCD) of the given numbers  150, 156, and 216.


You can decompose the numbers into product of primes

    150 = 2*3*5^2,

    156 = 2^2*3*13,

    216 = 2^3*3^3.


From these decompositions,  GCD(150, 156, 216) = 2*3 = 6.


So, the greatest possible cube has the edge of 6 centimeters long and the volume of 6^3 = 216 cm^3.

Solved.